OpenStudy (anonymous):

Assume the y varies inversely with x and y = 5 when x = 1.3. Write the indirect variation equation for the relationship.

OpenStudy (cwrw238):

y = k/x where k is a constant 5 = k / 1.3 k = 6.5 equation is y = 6.5 / x